An element X forms both a dichloride (XCl2) and a tetrachloride (XCl4). Treatment of 10.00g XCl2 with excess chloride forms 12.55 XCl4. Calculate the atomic mass of X, and identify X.
1
Expert's answer
2018-02-20T08:27:08-0500
Treatment of 10.00g XCl2 with excess chloride forms 12.55g XCl4. That means that the mass increases by 2.55g and it is the mass of 2 amounts of Cl. In XCl2 there are 2 amounts of Cl and 1 amount of X. In 10g XCl2 there are 2.55g of Cl and 10-2.55=7.45g of X (in such amount (ν) of XCl2).
If we take 1 mole of XCl2 there are 1 mole of X – A (g), where A – atomic mass, and 2 moles of Cl – 35.5×2=71 (g). The ratio between the masses of the elements is always the same :
2.55/7.45 = 71/A
A = 207.4 (g/mole) – Pb, lead.
Answer
Pb, 207.4 g/mole.